# Analytics, have some humility: a statistical view of fourth-down decision making

### 11 Dec 2023 v2

@@ 4 Results
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure ignores the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Thus in the previous Section 3 we modified the decision procedure to account for uncertainty. Our decision procedure is based on bootstrapping: judge our confidence in a decision by the proportion of bootstrapped win probability models that report the decision as optimal. Specifically, we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions.
In Section 3.3 we fit our fourth-down decision model, which given a fourth-down game-state outputs a ${\mathsf{boot\%}}$. In this Section we discuss our modeling results. We begin in Section 4.1 by illustrating our decision making process for example plays. Then in Section 4.2 we quantify the extent to which the proponents of the traditional decision procedure were overconfident in their fourth-down recommendations. We find that decision making changes substantially: far fewer fourth-down decisions are as obvious as analysts claim.
@@ 4.1 Example plays
We illustrate our decision procedure using example plays.
@@
Example play 1. First we compare Baldwin’s fourth-down decision making procedure to ours via example play 1 from Section 2.2. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal. In Figure 4 we incorporate uncertainty quantification into decision making. Although our point estimate (the blue column) suggests that ${\mathsf{Go}}$ provides a $1.4\%$ gain in win probability over ${\mathsf{FG}}$, our $90\%$ confidence interval $[-3.9\%,5.1\%]$ suggests that ${\mathsf{Go}}$ could either be a terrible or a great decision. Also, $50\%$ of the boostrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column), reflecting massive uncertainty in the optimal fourth-down decision. So, we are not confident in the fourth-down recommendation provided by the point estimate.
@@
Example play 2. Next we compare Burke’s fourth-down decision procedure to ours via example play 2 from Section 2.2. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot in Figure 2 (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary. That chart, however, says nothing about the estimated strength of making the optimal decision or about uncertainty quantification. Hence in Figure 5a we show our version of Burke’s chart, which uses color intensity to visualize the estimated gain in win probability by making a decision (darker colors indicate larger values).444 We use green for ${\mathsf{Go}}$, yellow for ${\mathsf{FG}}$, and red for ${\mathsf{Punt}}$ because we liked Burke’s Twitter thought that fourth-down decision making is like a traffic light. The pink dot (representing the actual play’s yardline and yards to go) lies in a moderately dark green region, indicating a moderate estimated gain in win probability by going for it. Being far from the decision boundary, however, does not imply it the best decision with certainty. Also, it is not obvious how far from the boundary is “far enough”. Hence in Figure 5b we provide an additional chart that illustrates uncertainty in the estimated optimal decision. Here, the color intensity indicates the proportion of bootstrapped models which make the estimated optimal decision. Much of the figure consists of dark colors, specifically dark green in the bottom and middle-left and dark red in the top and middle-right. For those combinations of yardline and yards to go, we are confident in the estimated optimal decision. For example play 2 (the pink dot), ${\mathsf{boot\%}}$ is high (about $90\%$), so we are confident that ${\mathsf{Go}}$ is the best decision. Other combinations of yardline and yards to go feature lighter colors, reflecting high uncertainty. For those game-states, we are not confident in the estimated optimal decision.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Example play 5. In Figure 8 we visualize our decision procedure for an infamous fourth-down play in which the Raiders had the ball against the Rams in Week 14 of 2022. ${\mathsf{Go}}$ provides a strong edge over ${\mathsf{Punt}}$ according to the ${\mathsf{WP}}$ point estimate ($+3.9\%$ ${\mathsf{WP}}$). Further, $92\%$ of the bootstrapped models say ${\mathsf{Go}}$ is better and our $90\%$ confidence interval of the estimated gain in win probability by going for it is $[-0.1\%,4.4\%]$, which is nearly positive. Thus, we are confident in this edge, and we strongly recommend that the Raiders should ${\mathsf{Go}}$555 In real life, the Raiders punted. Then, Rams quarterback Baker Mayfield countered with a successful 98 yard drive to win the game.
@@ 4.2 Analytics, have some humility
Since win probability point estimates are subject to substantial uncertainty, we should not simply recommend the decision that maximizes estimated win probability. Due to the random idiosyncracies in the history of football, which is contained in a relatively small dataset, win probability point estimates can make a decision seem strong even when it isn’t. In this Section we quantify the extent to which proponents of the traditional decision procedure are overconfident in their fourth-down recommendations.
In Figure 9 we compare traditional fourth-down recommendations to our recommendations. The strength of a traditional fourth-down decision recommendation is proportional to effect size, the estimated gain in win probability by making that decision ($x$-axis). The strength of our fourth-down recommendation is proportional to the bootstrap percentage. To faciliate easier communication to a non-technical football coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). Although the strength of a decision is proportional to the ${\mathsf{boot\%}}$, we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a decision recommendation. For all fourth-down plays from 2018-2022 in a given effect size bin, we plot the proportion of these decisions we view as confident, lean, and uncertain in Figure 9a. In Figure 9b we plot the frequency of these plays that lie in a given effect size bin.
@@
Finally, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. On this view, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 10 we rank coaches from 2018 to 2022 by the proportion of correctly made decisions that we are confident in. The top of this list has coaches like Kevin O’Connell and John Harbaugh from analytics bent organizations but also coaches that are considered more traditional on fourth down such as Ron Rivera, Sean McVay, and Mike Vrabel. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
@@ 5 Discussion
The standard mathematical approach to fourth-down decision making in American football is to make the decision that maximizes estimated win probability. The win probability estimates typically used today for fourth-down decision making arise from a statistical model fit from historical data. These machine learning models, however, are overfit high-variance estimators, exacerbated by the highly correlated nature of football play-by-play data. We conducted a simulation study to illuminate just how difficult it is to accurately fit a win probability model from historical data. Although these models find the general trend of true win probability, they are subject to substantial uncertainty. Existing decision procedures, based solely on effect size, ignore the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Hence we develop a machine learning framework that accounts for the correlated nature of observational football data and knits uncertainty quantification into our decision making. The practical football lesson arising from our new decision procedure is that far fewer fourth-down decisions are as obvious as analysts claim. Thus we ask football analysts to have some humility: for many game-states there is simply not enough data to be confident in win probability point estimates.
@@
In this study we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a fourth-down recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow that recommendation. Similarly, we believe it is arrogant for a football analyst to expect a coach to follow an uncertain fourth-down recommendation regardless of the effect size. In today’s age in which sports analysts are fighting a public relations war to be taken seriously by owners and coaches, overconfidence in decision recommendations doesn’t help non-believers buy what we’re selling. Further, the coach is an on-field expert who spends a significant amount of time with his players and has access to information that doesn’t show up in the data (i.e., unobserved confounders). For instance, Eagles coach Nick Sirianni may notice that his usually dominant offensive line is missing a key player today, which does not currently appear in the data. When the estimated optimal fourth-down decision has too much uncertainty, we suggest leaving the decision to the coach’s discretion. All that being said, although football analysts have been overconfident in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
In this work we used bootstrapping to drive uncertainty quantification. The way we tuned the bootstrap to acheive adequate marginal coverage in Section 3.2 is reminiscent of conformal inference. Conformal prediction intervals provide guarantees regarding the marginal coverage by ensuring every observation gets a holdout version of the interval (Lei et al.,, 2018). That the bootstrap methods with $\phi=1$ don’t acheive adequate coverage could be due to the time series dependence structure of football data. Recent work on conformal approaches for time series may help mitigate this issue (Angelopoulos et al.,, 2023). Nonetheless, valid distribution-free inference may not be possible in a binary outcome setting (Barber,, 2020). For instance, no confidence or prediction interval (e.g., $[0.63,0.71]$) will cover a 00 or $1$ outcome value.
Our analysis captures uncertainty in first-down win probability but not in decision transition probabilities (conversion probability, field goal success probability, and expected next yardline after punting). These models are fit from thousands of play-level i.i.d. observations, and so are subject to little sampling uncertainty, but are subject to nontrivial model uncertainty as they make simplifying assumptions. Conversion probability in particular is a delicate concept, as it depends on the offensive play call, the defensive play call, and the individual characteristics of each of the players on the field. A more fine-grained analysis would account for this additional uncertainty.
@@
Because statistical win probability models produce estimates that are too uncertain at many game-states, we suggest in future work exploring probabilistic state-space models to estimate win probability. Probabilistic models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and win probability is calculated by simulating games. The effective sample size of transition probability models is the number of plays because they are fit from independent play-level observations. Some analysts in private industry have created proprietary probabilistic win probability models, which they believe are more accurate than statistical models because they have a larger effective sample size. The tradeoff is that state-space models have higher bias than statistical models. Whereas both classes of models have hidden bias, or bias due to not adjusting for unobserved confounders, just state-space models have model bias as they make simplifying assumptions.
@@
Proprietors of probabilistic models believe that introducing bias in order to reduce variance improves the overall accuracy of the resulting win probability estimator. Nevertheless, these models are subject to their own set of issues, and we believe they aren’t as low-variance as some analysts claim. Properly modeling the distribution of the outcome of a play is nontrivial. In contrast to the simple binary win/loss outcome of statistical win probability models, the outcome variable of a transition probability model is the next game-state, which could include a change in yardline, score, or time. This distribution is quite complex: there is a spike at gaining 0 yards for incompletions, a spike for a touchdown, spikes for penalties, and other smooth possibly multimodal distributions for the outcome of run or pass plays, each of which change as a function of team quality and other confounders. Typical transition probability models are riddled with selection bias, as a coach generally chooses play calls that work for his specific players and don’t generalize to an “average” team. Further, uncertainty in these transition probabilities may, after being properly propagated through a state-space model, result in just as much (if not more) uncertainty in estimated win probability than estimates from statistical models. Additionally, one must take great care to carefully encode all the subtle rules of football into her model, and one needs sufficient computing power to simulate enough games to estimate win probability with enough granularity. We look forward to a public facing exploration of probabilistic win probability models in the future.
@@ Acknowledgments
Appendix A Estimating the win probability of a fourth-down decision
In this Section, we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability (denoted $\mathsf{WP_{1}}$) and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We estimate first-down win probability in Section 2.4 and decision transition probabilities in Appendix B.
@@ Appendix A Estimating the win probability of a fourth-down decision
$\sum_{y^{\prime}}(1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime}))\cdot{\mathbb{P}}(\text{yardline after punting is $y^{\prime}$}|{\mathbf{x}}).$ (A.1)
@@
${\mathbb{E}}_{\text{punt}}[1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime})|{\mathbf{x}}].$ (A.2)
@@
$1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{punt}}[y^{\prime}|{\mathbf{x}}],{\mathbf{x}}^{\prime}).$ (A.3)
We model the expected next yardline after punting as a function of yardline and punter quality in Appendix B.1.
@@
We model the probability of a successful field goal as a function of yardline and kicker quality in Appendix B.2. If the kicking team misses the field goal and the spot of the kick was within the 20 yardline, the opposing team has a first down at yardline $80$. If the kicking team misses the field goal and the spot of the kick was beyond the 20 yardline, the opposing team has a first down at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$ the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
@@
${\mathsf{WP}}(\text{make FG})=1-{\mathbb{E}}_{\text{kickoff}}[\mathsf{WP_{1}}(\text{yardline }y^{\prime},s^{\prime}=-s-3,{\mathbf{x}}^{\prime})].$ (A.6)
@@
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{kickoff}}[y^{\prime}],\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (A.7)
@@
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }75,\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (A.8)
@@
$\sum_{\Delta\geq z}\mathsf{WP_{1}}(\text{yardline }y-\Delta)\cdot{\mathbb{P}}(\text{gain $\Delta$ yards})+\sum_{\Delta<z}(1-\mathsf{WP_{1}}(\text{yardline }100-(y-\Delta)))\cdot{\mathbb{P}}(\text{gain $\Delta$ yards)}.$) . (A.9)
Here, the first-down win probability and conversion success probability are implicitly functions of game-state. As the probability distribution of the yards gained on a conversion attempt is complex, we employ a standard simplification from Burke, (2009). If the offensive team converts on fourth down and $z$ yards-to-go at yardline $y$, we assume they gain $z$ yards on that play, yielding a first down at yardline $y-z$ or a touchdown. If the offensive team fails to convert, we assume they turn the ball over at the initial yardline of the play, leaving the opposing team with a first down at yardline $100-y$. Thus the win probability of going for it is
$\begin{split}\mathsf{WP_{1}}(\text{yardline }y-z,{\mathbf{x}}^{\prime})\cdot{\mathbb{P}}(\text{convert}|{\mathbf{x}})+(1-\mathsf{WP_{1}}(\text{yardline }100-y,{\mathbf{x}}^{\prime}))\cdot(1-{\mathbb{P}}(\text{convert}|{\mathbf{x}})),\end{split}$ (A.10)
where ${\mathbf{x}}$ is the game-state on fourth-down and ${\mathbf{x}}^{\prime}$ is the game-state on the subsequent play (as before, ${\mathbf{x}}^{\prime}$ flips the game-state variables which are relative to the team with possession (e.g., score differential, team quality metrics, timeouts remaining) and doesn’t alter the other variables (e.g., time remaining)). We model conversion probability as a function of yardline, yards-to-go, and offensive and defensive quality in Appendix B.3.
Appendix B Estimating fourth-down decision transition probabilities
In this Section we estimate fourth-down decision transition probability models. In Section B.1 we model the expected next yardline after punting, in Section B.2 we model field goal success probability, and in Section B.3 we model conversion success probability.
B.1 Modeling punt outcome

### 19 Jan 2024 v4

@@ 4 Results
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure ignores the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Thus, we modified the decision procedure to account for uncertainty: we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Specifically, we measure our confidence in a decision by the proportion of bootstrapped win probability models that report the decision as optimal.
@@ 4.1 Example plays
We illustrate our decision procedure using example plays. Also, to compare our decision procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We detail the specification of this baseline coach model in Appendix H and include the model’s predictions in our decision figures.
Example play 1. First, we compare Baldwin’s fourth-down decision making procedure to ours via example play 1 from Section 2.2. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal, shown in Figure 1. In Figure 4 we incorporate uncertainty quantification into decision making. Although our point estimate (the blue column) suggests that ${\mathsf{Go}}$ provides a $1.4\%$ gain in win probability over ${\mathsf{FG}}$, our $90\%$ confidence interval $[-3.9\%,5.1\%]$ suggests that ${\mathsf{Go}}$ could either be a terrible or a great decision. Also, $50\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column), reflecting massive uncertainty in the optimal fourth-down decision. So, we are not confident in the fourth-down recommendation provided by the point estimate.
@@
Example play 2. Next, we compare Burke’s fourth-down decision procedure to ours via example play 2 from Section 2.2. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot in Figure 2 (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary. That chart, however, says nothing about the estimated strength of making the optimal decision or about uncertainty quantification. Hence, in Figure 5a we show our version of Burke’s chart, which uses color intensity to visualize the estimated gain in win probability by making a decision (darker colors indicate larger values).333 We use green for ${\mathsf{Go}}$, yellow for ${\mathsf{FG}}$, and red for ${\mathsf{Punt}}$ because we liked Burke’s Twitter thought that fourth-down decision making is like a traffic light. The pink dot (representing the actual play’s yardline and yards to go) lies in a moderately dark green region, indicating a moderate estimated gain in win probability by going for it. Being far from the decision boundary, however, does not imply it the best decision with certainty. Also, it is not obvious how far from the boundary is “far enough.” Hence in Figure 5b we provide an additional chart that illustrates uncertainty in the estimated optimal decision. Here, the color intensity indicates the proportion of bootstrapped models which make the estimated optimal decision. Much of the figure consists of dark colors, specifically dark green in the bottom and middle-left and dark red in the top and middle-right. For those combinations of yardline and yards to go, we are confident in the estimated optimal decision. For example play 2 (the pink dot), ${\mathsf{boot\%}}$ is high (about $90\%$), so we are confident that ${\mathsf{Go}}$ is the best decision. Other combinations of yardline and yards to go feature lighter colors, reflecting high uncertainty. For those game-states, we are not confident in the estimated optimal decision.
@@
We include more examples of our decision procedure in Appendix J. We also include an interactive Shiny App in which you can visualize our fourth-down decision procedure for any game-state on Github.444 https://github.com/snoopryan123/fourth_down/tree/main/3_shiny
@@ 4.2 Analytics, have some humility
Since win probability point estimates are subject to substantial uncertainty, we should not simply recommend the decision that maximizes estimated win probability. Due to the random idiosyncrasies in the history of football, which is contained in a relatively small dataset, win probability point estimates can make a decision seem strong even when it isn’t.
In Figure 6 we compare traditional fourth-down recommendations to our recommendations. The strength of a traditional fourth-down decision recommendation is proportional to effect size, the estimated gain in win probability by making that decision ($x$-axis). The strength of our fourth-down recommendation is proportional to the bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). For all fourth-down plays from 2018-2022 in a given effect size bin, we plot the proportion of these decisions we view as confident, lean, and uncertain in Figure 6a. In Figure 6b we plot the frequency of these plays that lie in a given effect size bin.
@@
Finally, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, in Appendix I we evaluate coaches only on plays for which we are confident.
@@ 5 Discussion
In this study we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a fourth-down recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow that recommendation. Similarly, we believe it is arrogant for a football analyst to expect a coach to follow an uncertain fourth-down recommendation, regardless of the effect size. In today’s age in which sports analysts are fighting a public relations war to be taken seriously by owners and coaches, overconfidence in decision recommendations doesn’t help non-believers buy what we’re selling. Further, the coach is an on-field expert who spends a significant amount of time with his players and has access to information that doesn’t show up in the data (i.e., unobserved confounders). For instance, Eagles coach Nick Sirianni may notice that his usually dominant offensive line is missing a key player today, which does not currently appear in the data. When the estimated optimal fourth-down decision has too much uncertainty, we suggest leaving the decision to the coach’s discretion. Having said that, although football analysts have been overconfident in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
Although our analysis improves the state of the art, it is not without limitations and there are many avenues for future work. We discuss these in detail below and in Appendix K.
@@
Because statistical win probability models produce estimates that are too uncertain at many game-states, we suggest that future work explore probabilistic state-space models to estimate win probability. Probabilistic models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and win probability is calculated by simulating games. The effective sample size of transition probability models is the number of plays because they are fit from independent play-level observations. Some analysts in private industry have created proprietary probabilistic win probability models, which they believe are more accurate than statistical models because they have a larger effective sample size. The tradeoff is that state-space models have higher bias than statistical models. Whereas both classes of models have hidden bias, or bias due to not adjusting for unobserved confounders, just state-space models have model bias as they make simplifying assumptions.
@@
Proprietors of probabilistic models believe that introducing bias in order to reduce variance improves the overall accuracy of the resulting win probability estimator. Nevertheless, these models are subject to their own set of issues, and we believe they aren’t as low-variance as some analysts claim. Properly modeling the distribution of the outcome of a play is nontrivial. In contrast to the simple binary win/loss outcome of statistical win probability models, the outcome variable of a transition probability model is the next game-state, which could include a change in yardline, score, time, or timeouts. This distribution is quite complex: there is a spike at gaining 0 yards for incompletions, a spike for a touchdown, spikes for penalties, and other smooth possibly multimodal distributions for the outcome of run or pass plays, each of which change as a function of team quality and other confounders. Typical transition probability models are riddled with selection bias, as a coach generally chooses play calls that work for his specific players and don’t generalize to an “average” team. Further, uncertainty in these transition probabilities may, after being properly propagated through a state-space model, result in just as much (if not more) uncertainty in estimated win probability than estimates from statistical models. Additionally, one must take great care to carefully encode all the subtle rules of football into her model, and one needs sufficient computing power to simulate enough games to estimate win probability with enough granularity. We look forward to a public facing exploration of probabilistic win probability models in the future.
@@ Acknowledgments
SUPPLEMENTARY MATERIAL
Appendix A Data details
variable variable description
win/loss 1 if the possession team wins the game, else 0
game seconds remaining num. seconds remaining in the game in $\{3600,...,1\}$
score differential
point differential between the offensive and defensive
team at the start of this play
total score total points scored during this game prior to this play
posteam spread pre-game point spread relative to the possession team
yardline
num. yards from the opponent’s endzone in
$\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
ydstogo
the number of yards the offense has left to gain
before achieving a first down or a touchdown
down a number in $\{1,2,3,4\}$ denoting the down of the play
posteam timeouts remaining num. timeouts in $\{3,2,1,0\}$ the offensive team has
defteam timeouts remaining num. timeouts in $\{3,2,1,0\}$ the defensive team has
receive 2h ko 1 if the possession team receives the $2^{nd}$ half kickoff, else 0
home 1 if the offensive team is at home, else 0
era
categorical variable grouping the year into
[…]
roof
categorical variable grouping the roof of the stadium into
[…]
game id unique identifier of this game
drive id unique identifier of this drive
posteam coach name of the coach of the possession team
Table 3: Game-state variables relevant to estimating win probability that describe the context at the start of a play.
Appendix B Estimating the win probability of a fourth-down decision
In this Section, we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability (denoted $\mathsf{WP_{1}}$) and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We estimate first-down win probability in Section 2.4 and decision transition probabilities in Appendix C.
@@ Appendix B Estimating the win probability of a fourth-down decision
$\sum_{y^{\prime}}(1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime}))\cdot{\mathbb{P}}(\text{yardline after punting is $y^{\prime}$}|{\mathbf{x}}).$ (11)
@@
${\mathbb{E}}_{\text{punt}}[1-\mathsf{WP_{1}}(\text{yardline }y^{\prime},{\mathbf{x}}^{\prime})|{\mathbf{x}}].$ (12)
@@
$1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{punt}}[y^{\prime}|{\mathbf{x}}],{\mathbf{x}}^{\prime}).$ (13)
We model the expected next yardline after punting as a function of yardline and punter quality in Appendix C.1.
@@
We model the probability of a successful field goal as a function of yardline and kicker quality in Appendix C.2. If the kicking team misses the field goal and the spot of the kick was within the 20 yardline, the opposing team has a first down at yardline $80$. If the kicking team misses the field goal and the spot of the kick was beyond the 20 yardline, the opposing team has a first down at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
@@
${\mathsf{WP}}(\text{make FG})=1-{\mathbb{E}}_{\text{kickoff}}[\mathsf{WP_{1}}(\text{yardline }y^{\prime},s^{\prime}=-s-3,{\mathbf{x}}^{\prime})].$ (16)
@@
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{kickoff}}[y^{\prime}],\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (17)
@@
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }75,\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (18)
@@
$\sum_{\Delta\geq z}\mathsf{WP_{1}}(\text{yardline }y-\Delta)\cdot{\mathbb{P}}(\text{gain $\Delta$ yards})+\sum_{\Delta<z}(1-\mathsf{WP_{1}}(\text{yardline }100-(y-\Delta)))\cdot{\mathbb{P}}(\text{gain $\Delta$ yards)}.$) . (19)
Here, the first-down win probability and conversion success probability are implicitly functions of game-state. As the probability distribution of the yards gained on a conversion attempt is complex, we employ a standard simplification from Burke, 2009b . If the offensive team converts on fourth down and $z$ yards-to-go at yardline $y$, we assume they gain $z$ yards on that play, yielding a first down at yardline $y-z$ or a touchdown. If the offensive team fails to convert, we assume they turn the ball over at the initial yardline of the play, leaving the opposing team with a first down at yardline $100-y$. Thus, the win probability of going for it is
$\begin{split}\mathsf{WP_{1}}(\text{yardline }y-z,{\mathbf{x}}^{\prime})\cdot{\mathbb{P}}(\text{convert}|{\mathbf{x}})+(1-\mathsf{WP_{1}}(\text{yardline }100-y,{\mathbf{x}}^{\prime}))\cdot(1-{\mathbb{P}}(\text{convert}|{\mathbf{x}})),\end{split}$ (20)
where ${\mathbf{x}}$ is the game-state on fourth down and ${\mathbf{x}}^{\prime}$ is the game-state on the subsequent play (as before, ${\mathbf{x}}^{\prime}$ flips the game-state variables which are relative to the team with possession (e.g., score differential, team quality metrics, timeouts remaining) and doesn’t alter the other variables (e.g., time remaining)). We model conversion probability as a function of yardline, yards-to-go, and offensive and defensive quality in Appendix C.3.
Appendix C Estimating fourth-down decision transition probabilities
In this Section we estimate fourth-down decision transition probability models. In Section C.1 we model the expected next yardline after punting, in Section C.2 we model field goal success probability, and in Section C.3 we model conversion success probability.
C.1 Modeling punt outcome